OpenStudy (anonymous):

decide whether the pair of lines is parallel, perpendicular, or neither 2x+4y=6 2x+4y=1

myininaya (myininaya):

write them both as y=mx+b

myininaya (myininaya):

if the slopes or opposite reciprocals of each other then they are perpendicular if the slopes are the same then they are parallel if neither of the above, then say neither

OpenStudy (anonymous):

They are parallel as the slopes are identical

OpenStudy (asadkarim7):

4y = -2x +6 4y = _ 2x +1 y = -1/2 x +6 y = -1/2 +1 m = -1/2 the lines are parallel